# Navigating Challenges in Spatio-temporal Modelling of Antarctic Krill Abundance: Addressing Zero-inflated Data and Misaligned Covariates

### v1

@@ Abstract
Antarctic krill (Euphausia superba) are among the most abundant species on our planet and serve as a vital food source for many marine predators in the Southern Ocean. In this paper, we utilise statistical spatio-temporal methods to combine data from various sources and resolutions, aiming to accurately model krill abundance. Our focus lies in fitting the model to a dataset comprising acoustic measurements of krill biomass. To achieve this, we integrate climate covariates obtained from satellite imagery and from drifting surface buoys (also known as drifters). Additionally, we use sparsely collected krill biomass data obtained from net fishing efforts (KRILLBASE) for validation. However, integrating these multiple heterogeneous data sources presents significant modelling challenges, including spatio-temporal misalignment and inflated zeros in the observed data. To address these challenges, we fit a Hurdle-Gamma model to jointly describe the occurrence of zeros and the krill biomass for the non-zero observations, while also accounting for misaligned and heterogeneous data sources, including drifters. Therefore, our work presents a comprehensive framework for analysing and predicting krill abundance in the Southern Ocean, leveraging information from various sources and formats. This is crucial due to the impact of krill fishing, as understanding their distribution is essential for informed management decisions and fishing regulations aimed at protecting the species.
Keywords: Antarctic krill; Hurdle model; Misaligned data; Spatial statistics; Zero-inflated.
@@ 1 Introduction
Integrating remotely sensed data and in situ measurements may provide a robust approach to addressing many challenges in modelling krill abundance. Satellite imagery offers large-scale, high-resolution information on key environmental variables (e.g., sea surface temperature, chlorophyll concentration, salinity, etc.), while in situ data provides precise, location-specific observations that capture dynamic oceanographic processes with finer detail. In this paper, the in situ covariates come from drifters, which track near-surface ocean currents and provide valuable insights into localised water movements. These drifter trajectories allow the derivation of additional environmental covariates, such as speed and mass flux, as we shall detail. When combined with satellite imagery, these datasets further enhance our ability to model the physical and biological factors influencing krill distribution. To estimate krill abundance, we use acoustic data (see Section 2.1.1), while the KRILLBASE dataset is employed for validation when extrapolating beyond the observed area.
@@ 2.1 Study Area and Sampling Approach
Figure 1: Study area (South Georgia, Subarea 48.3), showing acoustic data collected in 2016 and net haul data from KRILLBASE collected between 1926 and 2016. The green box indicates the region where acoustic data were collected, while the red box marks the area with sufficient net haul data to serve as a validation set for model predictions (see Section 4).
@@ 2.1.1 Acoustic Data
Acoustic surveys in South Georgia (Polar Ocean Ecosystem Time Series, Western Core Box) were conducted annually from 1997 to 2020, excluding the years 2002 and 2008, in intervals of 3 to 8 days within the December to February period (with December data considered as observations for the following year) (Fielding et al.,, 2014). Figure 4 (left column) presents the raw data for the first and last years, while plots for the remaining years are available in Figures SF2–SF21 (Supplementary Material). The surveys typically cover 8 transects, each 40 nautical miles in length, with a minimum separation of 10 nautical miles and a resolution of 500 metres.
The analysis of this dataset involves several challenges. Firstly, as shown in the Figure 4 (left column) and Figures SF2–SF21 (Supplementary Material), many observations are zero, indicating an absence of detected krill. Secondly, krill biomass can vary substantially even over short distances; in some instances, neighbouring observations span from zero to hundreds of $\text{g}/\text{m}^{2}$, highlighting the spatial heterogeneity of krill biomass distribution. Lastly, while we we would like to make predictions across the entire region shown in Figure 1, our data is limited to a much smaller area (green box, Figure 1). This limitation constrains our ability to generate reliable predictions for regions distant from the sampled locations. In Section 4, we present results from two analyses covering the regions outlined by the green and red boxes (Figure 1).
@@ 2.2.1 Satellite Imagery
Table 1 details all covariates used in our analysis, including products derived from satellite imagery, drifter trajectories (see Sections 2.2.2 and 3.2) and other environmental factors in the form of bathymetry (depth) and slope.
Table 1: Potential covariates for describing the spatial (and spatio-temporal) distribution of krill abundance. indicates covariates obtained from satellite imagery, and indicates covariates derived as products from drifter trajectories. § denotes covariates observed only during the months of December, January, and February (to match the acoustic data time window). indicates interpolation as described in Section SS1.1 (Supplementary Material).
@@
Speed (drifters) $0.01\times 0.01$—after interpolation 1997–2020§ Computed based on drifter trajectories speed_drif
@@ 3.1 Spatio-temporal Modelling
Let $\mathcal{X}\subset\mathbb{R}^{2}$ denote the continuous spatial domain, with observed locations $(s_{1},\cdots,s_{n})\subset\mathcal{X}$. Similarly, let $\mathcal{T}$ denote the temporal domain, where $t\in\mathcal{T}$. Following the notation in Krainski et al., (2018), let
@@
$\text{logit}(\pi_{it})=\beta_{0}^{z}+\beta_{1}^{z}\texttt{cov}_{1}^{z}+\cdots+\beta_{\ell_{1}}^{z}\texttt{cov}_{\ell_{1}}^{z}+\psi_{it},$ (1)
@@
$\log(\mu_{it})=\beta_{0}^{y}+\beta_{1}^{y}\texttt{cov}_{1}^{y}+\cdots+\beta_{\ell_{2}}^{y}\texttt{cov}_{\ell_{2}}^{y}+\gamma\cdot\psi_{it}+\xi_{it},$ (2)
where $\ell_{1}$ and $\ell_{2}$ denote the number of covariates in each model, which may overlap, and $\psi_{it}$ and $\xi_{it}$ are spatio-temporal random effects. Here, $\gamma$ serves as a “copy” factor to scale the shared random effect $\psi_{it}$ in the biomass model, allowing for dependencies between the presence-absence and biomass components.
@@
$\psi_{it}=\alpha\psi_{i,(t-1)}+\phi_{it},$ (3)
where $|\alpha|<1$, $\psi_{i1}\sim\text{Normal}(0,\sigma^{2}_{\phi}/(1-\alpha^{2}))$, and $\phi_{it}$ is a temporally independent but spatially dependent Gaussian Process (GP) at each year with covariance given by a Matérn kernel, i.e.,
$\text{Cov}(\phi_{it},\phi_{jt})=\frac{\sigma^{2}}{2^{\nu-1}\Gamma(\nu)}(\kappa\cdot h)^{\nu}K_{\nu}(\kappa\cdot h),$ (4)
where $h=||s_{it}-s_{jt}||$ is the Euclidean distance between the locations $s_{it}$ and $s_{jt}$, and $\sigma^{2}$ denotes the marginal variance. $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ is a modified Bessel function of the second kind, such that $\nu>0$ determines the mean square differentiability of the corresponding process. Lastly, $\kappa>0$ is related to the range $\rho$, such that $\rho=\sqrt{8\nu}/\kappa$.
@@ 3.1.1 Spatial Modelling
In Section 4.2, we will perform an aggregated spatial analysis and, in this instance, drop the temporal component from Equations (1) and (2). In this setting, the corresponding spatial Hurdle-Gamma model will be defined as before; however, the random effects $\psi_{i}$ and $\nu_{i}$ will be modelled as Gaussian processes in space only, using a Matérn kernel similar to that in Equation (4).
@@ 3.2 Deriving Products from Drifter Trajectories
  • Speed (drifters): we compute the speed of all drifter observations given by $|z_{i}(t)|$ and then map these to their corresponding locations ${q}_{i}(t)$. After which we create a spatially gridded product at the desired resolution by interpolating using Gaussian processes with a Matérn kernel, as detailed in Section SS1.1 (Supplementary Material). Note, importantly, that this covariate is expected to be different from the speed (satellite) covariate in Table 1, as the satellite data we use provides estimates of the geostrophic velocity computed from sea surface height (SSH) gradients, whereas drifter speeds are expected to be a mix of geostrophic and ageostrophic velocities (caused for example by surface winds). This difference is explained in detail by O’Malley et al., (2023). Figure 3 shows maps of the speed estimates in Subarea 48.3 from satellite data and drifter observations for comparison.
@@
    where $\Delta$ is the temporal sampling interval and is assumed constant (which it is with the drifter data used in this paper; see Section 2.2.2). The sequence $\{h_{j}\}$ in Equation (6) is known as a data taper that satisfies $\sum_{j=1}^{n_{i}}h_{j}^{2}=1$, where we select $\{h_{j}\}$ to be a DPSS (discrete prolate spheroidal sequence) of order 1 (with bandwidth parameter set to 4), and is used to remove bias in the estimate of the spectrum, see (Percival and Walden,, 1993, Chapter 6) for more details. We can then approximate the expected frequency in Equation (5) by
@@
    are the observed Fourier frequencies and $\kappa=\sum_{k=1}^{n_{i}}\hat{S}_{i}(\omega_{k})$. The calculation of expected frequency requires the velocity time series to be approximately stationary, which will not generally be the case for an entire drifter trajectory in our region of interest. Therefore, we compute the expected frequency for each drifter trajectory over temporal windows (with 50% overlap) of length 5 days, which is considered to be a good approximation of the “decorrelation timescale” (i.e., the timescale at which a drifter “forgets” its history of movement), and is a standard choice in ocean drifter analysis (see O’Malley et al., (2021) and references therein). Finally, we derive a spatial gridded map by taking the set of computed expected frequencies and mapping them onto the midpoint location of each trajectory segment and then spatially smoothing onto a grid using Gaussian processes, as detailed in Section SS1.1 (Supplementary Material).
@@
  • Mass flux: the residence time computes the average time a drifter continuously spends in a fixed spatial region. A natural orthogonal covariate to also include is the number of drifters that pass through this region over time. This can be interpreted as the mass flux of drifters as it measures the rate at which drifters move across a unit area per unit of time. The motivation to include this covariate also comes from Murphy et al., (2004), who find associations between water volume flux (which the drifters are mimicking near the surface) and krill flux. For our analysis, mass flux is computed in exactly the same way as residence time, see Section SS1 (Supplementary Material) for details.
  • Density of drifters: lastly, as the drifters are freely floating then they are not uniformly sampling the ocean and instead are likely to be preferentially sampling the ocean due to the impact of, for example, convergent or divergent zones (Middleton and Garrett,, 1986). Although the density of drifters and krill will not necessarily aggregate in the same way, it nonetheless could be informative as a covariate. We therefore include a basic estimate of the density of drifters which corresponds to the total number of hours spent by drifters in each pixel of the spatial image, as detailed in Section SS1 (Supplementary Material).
@@
Figure 3: Covariates in South Georgia (Subarea 48.3), as described in Table 1. All drifter-derived products were computed based on the trajectories collected from 1997 to 2020. 1: Speed ($\text{m}\cdot\text{s}^{-1}$) from the satellite, averaged over 1997–2020. 2: Speed ($\text{m}\cdot\text{s}^{-1}$) from the drifters. 3: Expected frequency ($\text{day}^{-1}$). 4: Residence time ($\text{min}\cdot\text{km}^{-2}$). 5: Mass flux ($\text{segments}\cdot\text{t}^{-1}\cdot\text{km}^{-2}$), where “time” refers to the entire observational period, i.e., 24 years. 6: Density of drifters ($\text{buoys}\cdot\text{km}^{-2}$).
@@ 4 Results
Following the modelling framework described in Section 3.1 and incorporating covariates from satellite imagery and drifter products (Table 1), we fit a Hurdle-Gamma model for krill biomass from acoustic data (Section 2.1.1) under two settings. First, we apply the model to the disaggregated data at its original spatio-temporal resolutions and focus on the region where we observed the data (green box, Figure 1). In this setting, our primary interest lies in the interpretability of some model parameters. Second, to enhance predictive capability outside the observed window (red box, Figure 1) and in line with the approach of Warwick-Evans et al., (2022), we fit a spatial-only version of our model to the acoustic data aggregated across space and time. This setup also enables variable selection at a feasible computational cost.

### v2

@@ Abstract
Antarctic krill (Euphausia superba) are among the most abundant species on our planet and serve as a vital food source for many marine predators in the Southern Ocean. In this paper, we utilise statistical spatio-temporal methods to combine data from various sources and resolutions, aiming to model krill abundance. Our focus lies in fitting the model to a dataset comprising acoustic measurements of krill biomass. To achieve this, we integrate climate covariates obtained from satellite imagery and from drifting surface buoys (also known as drifters). Additionally, we use sparsely collected krill biomass data obtained from net fishing efforts (KRILLBASE) for validation. However, integrating these multiple heterogeneous data sources presents significant modelling challenges, including spatio-temporal misalignment and inflated zeros in the observed data. To address these challenges, we fit a Hurdle-Gamma model to jointly describe the occurrence of zeros and the krill biomass for the non-zero observations, while also accounting for misaligned and heterogeneous data sources, including drifters. Therefore, our work presents a comprehensive framework for analysing and predicting krill abundance in the Southern Ocean, leveraging information from various sources and formats. This is crucial due to the impact of krill fishing, as understanding their distribution is essential for informed management decisions and fishing regulations aimed at protecting the species.
Keywords: Antarctic krill; Hurdle model; Misaligned data; Zero-inflated.
@@ 1 Introduction
Integrating remotely sensed data and in situ measurements may provide a robust approach to addressing many challenges in modelling krill abundance. Satellite imagery offers large-scale, high-resolution information on key environmental variables (e.g., sea surface temperature, chlorophyll concentration, salinity, etc.), while in situ data provides precise, location-specific observations that capture dynamic oceanographic processes with finer detail. In this paper, the in situ covariates come from drifters, which track near-surface ocean currents and provide valuable insights into localised water movements. These drifter trajectories allow the derivation of additional environmental covariates, such as surface speed and mass flux, as we shall detail. When combined with satellite imagery, these datasets further enhance our ability to model the physical and biological factors influencing krill distribution. To estimate krill abundance, we rely exclusively on the acoustic observations (see Section 2.1.1), while the KRILLBASE dataset is employed for validation when extrapolating beyond the observed area. Our analysis proceeds in two directions: (I) a disaggregated spatio-temporal model that exploits the exact locations of the acoustic records, and (II) a spatial model fitted to spatio-temporally aggregated observations, aimed at predicting krill abundance across a wider area—each setting offers a distinct yet coherent view of krill distribution.
@@ 2.1 Study Area and Sampling Approach
Figure 1: Study area (South Georgia, Subarea 48.3), showing sampling locations of acoustic data collected in 2016 and net haul data from KRILLBASE collected between 1926 and 2016. The green box indicates the region where acoustic data were collected, while the red box marks an area selected through visual assessment, where the net haul data were considered to offer a useful basis for comparison with model predictions (see Section 4).
@@ 2.1.1 Acoustic Data
Acoustic surveys in South Georgia (Polar Ocean Ecosystem Time Series, Western Core Box) were conducted annually from 1997 to 2020, excluding the years 2002 and 2008, in intervals of 3 to 8 days within the December to February period (with December data considered as observations for the following year) (Fielding et al.,, 2014). Figure 4 (left column) presents the raw data for the first and last years, while plots for the remaining years are available in Figures SF4–SF23 (Supplementary Material). The surveys typically cover 8 transects, each 40 nautical miles in length, with a minimum separation of 10 nautical miles and a resolution of 500 metres.
The analysis of this dataset involves several challenges. Firstly, as shown in the Figure 4 (left column) and Figures SF4–SF23 (Supplementary Material), many observations are zero, indicating an absence of detected krill. Secondly, krill biomass can vary substantially even over short distances; in some instances, neighbouring observations span from zero to hundreds of $\text{g}/\text{m}^{2}$, highlighting the spatial heterogeneity of krill biomass distribution. Lastly, while we would like to make predictions across the entire region shown in Figure 1, our data is limited to a much smaller area (green box, Figure 1). This limitation constrains our ability to generate reliable predictions for regions distant from the sampled locations. In Section 4, we present results from two analyses covering the regions outlined by the green and red boxes (Figure 1).
@@ 2.2.1 Satellite Imagery
Table 1 lists all covariates used in our analysis, including products derived from satellite imagery, drifter trajectories (see Sections 2.2.2 and 3.2) and other environmental factors in the form of bathymetry (depth) and slope (calculated from the bathymetry). Section SS1.1 (Supplementary Material) provides a brief description of the satellite imagery from the Copernicus Marine Service used.
Table 1: Potential covariates for describing the spatial (and spatio-temporal) distribution of krill abundance. indicates covariates obtained from satellite imagery, and indicates covariates derived as products from drifter trajectories. § denotes covariates observed only during the months of December, January, and February (to match the acoustic data time window). indicates interpolation as described in Section SS1.2.1 (Supplementary Material).
@@
Surface speed (drifters) $0.01\times 0.01$—after interpolation 1997–2020§ Computed based on drifter trajectories speed_drif
@@ 3.1 Spatio-temporal Modelling
Let $\mathcal{X}\subset\mathbb{R}^{2}$ denote the continuous spatial domain, with observed locations $(s_{1},\cdots,s_{n})\subset\mathcal{X}$. Similarly, we define $\mathcal{T}$ as the temporal domain, with $t\in\{1,\cdots,T\}$ indexing discrete time points. Following the notation in Krainski et al., (2018), let
@@
$\text{logit}(\pi_{it})=\beta_{0}^{z}+\beta_{1}^{z}\texttt{cov}_{1,it}^{z}+\cdots+\beta_{\ell_{1}}^{z}\texttt{cov}_{\ell_{1},it}^{z}+\psi_{it},$ (1)
@@
$\log(\mu_{it})=\beta_{0}^{y}+\beta_{1}^{y}\texttt{cov}_{1,it}^{y}+\cdots+\beta_{\ell_{2}}^{y}\texttt{cov}_{\ell_{2},it}^{y}+\gamma\cdot\psi_{it}+\xi_{it},$ (2)
where $\ell_{1}$ and $\ell_{2}$ denote the number of covariates in each model, which may overlap, and $\psi_{it}$ and $\xi_{it}$ are spatio-temporal random effects. Here, $\gamma$ serves as a “copy” factor to scale the shared random effect $\psi_{it}$ in the biomass model, allowing for dependencies between the presence-absence and biomass components. We note that, although the covariate effects are specified linearly in Equations (1) and (2), they could be replaced by mildly non-linear bases—such as cubic B-splines with a small knot set (Fahrmeir et al.,, 2004)—if future analyses indicate a clear benefit.
@@
$\psi_{it}=\alpha_{\psi}\psi_{i,(t-1)}+\phi_{it},$ (3)
where $|\alpha_{\psi}|<1$, $\psi_{i1}\sim\text{Normal}(0,\sigma^{2}_{\phi}/(1-\alpha_{\psi}^{2}))$, and $\phi_{it}$ is a temporally independent but spatially dependent Gaussian Process (GP) at each year with covariance given by a Matérn kernel, i.e.,
$\text{Cov}(\phi_{it},\phi_{jt})=\frac{\sigma^{2}_{\phi}}{2^{\nu-1}\Gamma(\nu)}(\kappa\cdot h)^{\nu}K_{\nu}(\kappa\cdot h),$ (4)
where $h=||s_{it}-s_{jt}||$ is the Euclidean distance between the locations $s_{it}$ and $s_{jt}$, and $\sigma^{2}_{\phi}$ denotes the marginal variance. $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ is a modified Bessel function of the second kind, such that $\nu>0$ determines the mean square differentiability of the corresponding process. Lastly, $\kappa>0$ is related to the range $\rho$, such that $\rho=\sqrt{8\nu}/\kappa$.
@@ 3.1.1 Spatial Modelling
In Section 4.2, we perform an aggregated spatial analysis and, in this instance, drop the temporal component from Equations (1) and (2). In this setting, the corresponding spatial Hurdle-Gamma model will be defined as before; however, the random effects $\psi_{i}$ and $\nu_{i}$ will be modelled as Gaussian processes in space only, using a Matérn kernel similar to that in Equation (4).
This spatial-only formulation can be interpreted as an empirical approximation to the mean (averaged over time) of the full spatio-temporal model; i.e., after averaging the original response $z_{it}\cdot y_{it}$ over the survey years, its model-based expectation $(1/T)\sum_{t}\pi_{it}\cdot\mu_{it}$ can be approximated by the product of the time-averaged components $\bar{\pi}_{i}\cdot\bar{\mu}_{i}$. That simplification relies on three working conditions—(I) covariate effects do not vary with year, (II) no strong residual year trend remains once covariates are included, and (III) the underlying biomass field shows no long-term drift—under which temporally averaged covariates and a purely spatial random effect provide a coherent large-scale picture of krill distribution. However, if any of these conditions is violated, a model that retains an explicit time dimension would be preferable. In Section SS2 (Supplementary Material), we illustrate this equivalence by means of a simulation study.
@@ 3.2 Deriving Products from Drifter Trajectories
  • Surface speed (drifters): we compute the speed of all drifter observations given by $|z_{i}(t)|$ and then map these to their corresponding locations ${q}_{i}(t)$. After which we create a spatially gridded product at the desired resolution by interpolating using Gaussian processes with a Matérn kernel, as detailed in Section SS1.2.1 (Supplementary Material). Note, importantly, that this covariate is expected to be different from the speed (satellite) covariate in Table 1, as the satellite data we use provides estimates of the geostrophic velocity computed from sea surface height (SSH) gradients, and is averaged over different depths, whereas drifter speeds are expected to be a mix of geostrophic and ageostrophic velocities (caused for example by surface winds) at or near the surface. This difference is explained in detail by O’Malley et al., (2023). Figure 3 shows maps of the surface speed estimates in Subarea 48.3 from satellite data and drifter observations for comparison.
@@
    where $\Delta$ is the temporal sampling interval and is assumed constant (which it is with the drifter data used in this paper; see Section 2.2.2). The sequence $\{h_{j}\}$ in Equation (6) is known as a data taper that satisfies $\sum_{j=1}^{n_{i}}h_{j}^{2}=1$, where we select $\{h_{j}\}$ to be a DPSS (discrete prolate spheroidal sequence) of order 1 (with bandwidth parameter set to 4), and is used to remove bias in the estimate of the spectrum, see Percival and Walden, (1993, Chapter 6) for more details. We can then approximate the expected frequency in Equation (5) by
@@
    are the observed Fourier frequencies and $\kappa=\sum_{k=1}^{n_{i}}\hat{S}_{i}(\omega_{k})$. The calculation of expected frequency requires the velocity time series to be approximately stationary, which will not generally be the case for an entire drifter trajectory in our region of interest. Therefore, we compute the expected frequency for each drifter trajectory over temporal windows (with 50% overlap) of length 5 days, which is considered to be a good approximation of the “decorrelation timescale” (i.e., the timescale at which a drifter “forgets” its history of movement), and is a standard choice in ocean drifter analysis (see O’Malley et al., (2021) and references therein). Finally, we derive a spatial gridded map by taking the set of computed expected frequencies and mapping them onto the midpoint location of each trajectory segment and then spatially smoothing onto a grid using Gaussian processes, as detailed in Section SS1.2.1 (Supplementary Material).
@@
  • Mass flux: the residence time computes the average time a drifter continuously spends in a fixed spatial region. A natural orthogonal covariate to also include is the number of drifters that pass through this region over time. This can be interpreted as the mass flux of drifters as it measures the rate at which drifters move across a unit area per unit of time. The motivation to include this covariate also comes from Murphy et al., (2004), who find associations between water volume flux (which the drifters are mimicking near the surface) and krill flux. For our analysis, mass flux is computed in exactly the same way as residence time, see Section SS1.2 (Supplementary Material) for details.
  • Density of drifters: lastly, as the drifters are freely floating then they are not uniformly sampling the ocean and instead are likely to be preferentially sampling the ocean due to the impact of, for example, convergent or divergent zones (Middleton and Garrett,, 1986). Although the density of drifters and krill will not necessarily aggregate in the same way, it nonetheless could be informative as a covariate. We therefore include a basic estimate of the density of drifters which corresponds to the total number of hours spent by drifters in each pixel of the spatial image, as detailed in Section SS1.2 (Supplementary Material).
@@
Figure 3: Covariates in South Georgia (Subarea 48.3), as described in Table 1. All drifter-derived products were computed based on the trajectories collected from 1997 to 2020. 1: Speed ($\text{m}\cdot\text{s}^{-1}$) from the satellite, averaged over 1997–2020. 2: Surface speed ($\text{m}\cdot\text{s}^{-1}$) from the drifters. 3: Expected frequency ($\text{day}^{-1}$). 4: Residence time ($\text{min}\cdot\text{km}^{-2}$). 5: Mass flux ($\text{segments}\cdot\text{t}^{-1}\cdot\text{km}^{-2}$), where “time” refers to the entire observational period, i.e., 24 years. 6: Density of drifters ($\text{buoys}\cdot\text{km}^{-2}$).
@@ 4 Results
Following the modelling framework described in Section 3.1 and incorporating covariates from satellite imagery and drifter products (Table 1), we fit a Hurdle-Gamma model for krill biomass from acoustic data only (Section 2.1.1) under two settings. First, we apply the model to the disaggregated data at its original spatio-temporal resolutions and focus on the region where we observed the data (green box, Figure 1). In this setting, our primary interest lies in the interpretability of some model parameters. Second, to enhance predictive capability outside the observed window (red box, Figure 1) and in line with the approach of Warwick-Evans et al., (2022), we fit a spatial-only version of our model to the acoustic data aggregated across space and time. This setup also enables variable selection at a feasible computational cost.